OpenStudy (anonymous):

A bag contains four red counters and six black counters. A counter is picked at random and not replaced. A second counter is then picked. Find the probability that the second counter is red, given that the first counter is red.

OpenStudy (anonymous):

a dice is in the form of a tetrahedron and its faces are marked 1,2,3 and 4. the score is the number on which the dice lands, if two tetrahedral dice are thrown find the probability that the sum of the two scores is 5

OpenStudy (anonymous):

for the first one we know one red counter is picked first how many red counters are left?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

? there are 4 to begin with, then one red counter is picked how may red counters are left after one is picked?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

right and how may counters are left in the bag?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

yes so the probability that the second counter is red, given that the first one is, is \(\frac{3}{9}=\frac{1}{3}\)

OpenStudy (anonymous):

great thank u and for the second one?

OpenStudy (anonymous):

you have two dice, each with 4 sides, so the total number of possible outcomes is \(4\times 4=16\) now we have to figure out how many ways there are to get a total of \(5\)

OpenStudy (anonymous):

i got 4 ways

OpenStudy (anonymous):

yeah me too

OpenStudy (anonymous):

so the probability you are looking for is \(\frac{4}{16}=\frac{1}{4}\)

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

but is there a formula for the probability of the dice